You take out a $7,400 car loan that calls for 36 monthly payments starting after 1 month at an APR of 9%.

a. What is your monthly payment? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

- monthly payment?

b. What is the effective annual interest rate on the loan? (Do not round intermediate calculations. Enter your answer as a percent

-effective annual interest rate??

c. Now assume the payments are made in four annual year-end installments. What annual payment would have the same present value as the monthly payment you calculated? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

what is annal paynment?

​​​​​​​
Answer #1

a.Information provided:

Present value= $7,400

Time= 36 months

Interest rate= 9%/12= 0.75%

The amount of monthly payment is calculated by entering the below in a financial calculator:

PV= -7,400

N= 36

I/Y= 0.75

Press the CPT key and PMT to compute the amount of monthly payment.

The value obtained is 235.3180.

Therefore, the amount of monthly payment is $235.32.

b. Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.09/12)^12-1

= 1.0938-1

= 0.0938*100= 9.3807%

Therefore, the effective annual rate is 9.38%.

c.The amount of annual payment is calculated by entering the below in a financial calculator:

PV= -7,400

N= 4

I/Y= 9

Press the CPT key and PMT to compute the amount of annual payment.

The value obtained is 2,284.15.

Therefore, the amount of annual payment is $2,284.15.

In case of any query, kindly comment on the solution